# Understanding and Estimating the Execution Time of Quantum Programs

### v1

@@ Abstract
Due to the scarcity of quantum computing resources, researchers and developers have very limited access to real quantum computers. Therefore, judicious planning and utilization of quantum computer runtime are essential to ensure smooth execution and completion of projects. Accurate estimation of a quantum program’s execution time is thus necessary to prevent unexpectedly exceeding the anticipated runtime or the maximum capacity of the quantum computers; it also allows quantum computing platforms to make precisely informed provisioning and prioritization of quantum computing jobs. In this paper, we first study the characteristics of quantum programs’ runtime on simulators and real quantum computers. Then, we introduce an innovative method that employs a graph transformer-based model, utilizing the graph information and global information of quantum programs to estimate their execution time. We selected a benchmark dataset comprising over 1510 quantum programs, initially predicting their execution times on simulators, which yielded promising results with an R-squared value over 95%. Subsequently, for the estimation of execution times on quantum computers, we applied active learning to select 340 samples with a confidence level of 95% to build and evaluate our approach, achieving an average R-squared value exceeding 90%. Our approach can be integrated into quantum computing platforms to provide an accurate estimation of quantum execution time and be used as a reference for prioritizing quantum execution jobs. In addition, our findings provide insights for quantum program developers to optimize their programs in terms of execution time consumption, for example, by prioritizing one-qubit gates over two-qubit gates.
@@
Keywords: Quantum Computing, Quantum Circuits, Quantum Computers, Quantum Execution Time, Graph Transformer
1 Introduction
@@
Over the last decade, there has been a significant transformation in quantum computing, evolving from experimental physics to cloud-accessible hardware (e.g., the IBM Quantum or Amazon Braket platforms), indicating a shift towards early-stage commercialization. This evolution has engaged a diverse audience, leading to research into practical quantum algorithms and gaining attention from academia, governments, and businesses. Notably, recent demonstrations of quantum utility [12] suggest that quantum computers may eventually outperform classical counterparts for real-world tasks [38].
@@
In quantum computing, the process of making a quantum program (a.k.a, quantum circuit) circuit executable on a specific backend is akin to classical computing’s transformation of high-level code into low-level, machine-specific code. Designed initially without considering hardware constraints such as gate limitations or connectivity issues, a quantum circuit undergoes a series of compilation passes [18]. These passes progressively adapt the circuit to the target backend’s restrictions. Given the complexity of these compilation problems, there is a variety of available approaches in various quantum SDKs and tools. These SDKs differ in the programming languages, gate sets, and quantum computer vendors they support [22]. For example, vendor-agnostic SDKs such as pytket [25] accommodate various languages and gate sets, whereas vendor-specific SDKs like IBM’s Qiskit [9] are restricted to their proprietary systems.
@@
Executing a quantum circuit on a quantum computer involves exclusively using the quantum computer for a specific period, similar to the batch processing mode used in old mainframe systems. Users submit their designed and compiled quantum circuit to a pre-selected quantum computer, wait in a queue for processing, and finally the quantum computer executes the quantum circuit. The time taken to execute the quantum circuit is referred to as the “execution time”. Given the scarcity and the substantial cost associated with quantum computing resources, the window available for using these state-of-the-art machines is exceptionally limited. Using IBM’s Quantum Platform  [^1] https://quantum.ibm.com/ as an example, free users are limited to accessing two to three quantum computers per month, and usage is restricted to no more than 10 minutes. Any additional usage time is billed by the second and can be quite costly (\$1.6 USD per second at the time of writing). Exceeding time constraints can lead to incomplete executions or cause significant costs, potentially delaying critical projects or incurring other severe consequences. Prudent planning and strategic utilization of quantum computer time are imperative to ensure the smooth execution and completion of projects. This underscores the necessity for an accurate estimation of quantum circuits’ execution time. In addition, an accurate estimation of the execution time can help quantum computing platform providers (e.g., IBM Quantum Platform) with their provisioning and prioritization of quantum computing jobs.
@@
Despite the crucial need for reliable execution time predictions for quantum programs, current research in this area remains sparse. The IBM Quantum Platform’s provided estimates carry substantial inaccuracies, failing to meet the demands of practitioners and researchers in the quantum computing field adequately. There are existing studies on estimating the execution time of classical programs [7, 10, 13, 2]. For instance, Huang et al. [7] collected a set of program execution features of a classical program, such as loop counts (the number of times a particular loop executed) and branch counts (the number of times a particular conditional branch executed), and build polynomial regression models to predict its execution time. This method achieved an accuracy of over 93% in predicting the execution time for the programs they tested. However, these approaches for classical programs do not apply to quantum programs, for three reasons: 1) quantum programs and classical programs are different in nature: classical programs describe the execution flows of code statements, whereas quantum programs describe the structure of a quantum circuit; 2) most approaches [7, 10, 13] require tracing runtime information from the subject programs to collect features for model training or inference, whereas tracing quantum programs is not possible, at least for now, due to the famous no-cloning theorem [35]; 3) these approaches build a model for each program [7, 10, 2] or similar programs [13] as different programs have very different features (e.g., different loops or branches), but it is too expensive to collect sufficient training data for each quantum program or similar ones, as quantum computing resources are rather limited.
@@
To address this gap, this paper introduces an innovative methodology that harnesses graph transformer-based models to predict the execution times of quantum circuits, leveraging the topological graph data intrinsic to these circuits and their global characteristics. We have analyzed over 1510 quantum circuits, ranging in size from 2 to 127 qubits, and executed them using different backends: two simulators and two real quantum computers. We organize our study by answering three research questions.
@@
    Understanding the execution times of quantum circuits is essential for effectively managing the utilization of quantum computing resources. Besides, it is only meaningful to forecast the execution time of a specific quantum circuit on a specific quantum backend when the execution times of different quantum circuits vary on different quantum backends. We observe that the execution times of different quantum circuits are noticeably different, whether on simulators or real quantum computers. Furthermore, the IBM Quantum Platform’s estimation of execution times for quantum circuits on quantum computers is not sufficient, which leads us to the second and third research questions.
@@
    Quantum program developers typically use simulators to verify their programs before executing them on real quantum computers. Time-consuming simulation can cost significant computing resources (e.g., IBM Quantum’s cloud-based simulators) and impact the productivity of quantum program development. Thus, understanding the simulation time of quantum programs can help quantum developers better plan their simulation jobs (e.g., determining running a program for more or fewer shots) and help simulation service providers optimize the management of the simulation jobs. In this RQ, we utilize various features, including features that capture the overall characteristics of a circuit (i.e., global features and features that capture the graph structure of a circuit (i.e., graph features), and construct machine learning models to estimate the execution time of quantum programs on simulators. Our results demonstrate the precision of our automated estimation and highlight the influential features.
@@
    Quantum computing resources are very limited currently and in the foreseeable future. Understanding the resource requirement (in terms of execution time) of quantum circuits before they execute is important for the optimal planning and allocation of quantum computing resources. Thus, this research question investigates the feasibility of creating a model that can accurately predict the execution times of quantum programs on real quantum computers. Facing the limited quantum resources, we use active learning to collect the most representative training data. Our results show that we can achieve accurate predictions with a relatively small amount of data, substantially outperforming the estimation provided by the IBM Quantum Platform, and highlight the difference between the important predictors for the simulation time prediction and the prediction for real computer execution time.
@@
  • We provided an in-depth understanding of the execution time of quantum programs on both simulators and real quantum computers.
@@
  • We identified the influential features that explain the quantum program execution time, providing insights for quantum program developers to consider when developing efficient quantum programs.
@@
Paper organization. Section 2 discussed prior work related to our study. Section 3 introduces the background knowledge of quantum computing, including quantum circuits, compilation, and execution. Section 4 describes the data collection and preparation process. Section 5 presents our model construction and the evaluation methods. Section 6 presents and discusses our results for answering the research questions. Section 7 discusses the limitations of our study. Finally, Section 8 concludes our study and suggests directions for future research.
2 Related Works
@@
As far as we know, there exist no prior work on understanding or estimating the execution time of quantum programs. In this section, we discuss three categories of prior works that are closest to ours: quantum resource estimation, quantum program fidelity estimation, and classical software execution time estimation.
2.1 Quantum resource estimation
@@
Quantum resource estimation offers a forward-looking solution for implementing key quantum algorithms by predicting necessary resources—such as qubits and runtime—based on future hardware projections, without running the actual programs. This approach allows consideration of quantum programs beyond current simulator and computer capabilities [21]. QuRE [27] was among the pioneering efforts on quantum resource estimation. The authors developed a resource estimation methodology for diverse quantum technologies and algorithms, capable of automatically estimating metrics like qubit count, completion time, success probability, and gate counts for different physical gate types. Later on, another work has been undertaken to develop a framework for quantum resource estimation and evaluate the resource requirements for large-scale quantum applications, with a particular emphasis on the critical role of qubit parameters, aiming to expedite progress towards achieving practical quantum advantage [1]. Another recent research provided resource estimates for quantum derivative pricing; it introduced a resource-efficient re-parameterization method, offered insights into resource demands for benchmark use cases, and outlined a roadmap for future quantum algorithm and hardware enhancements [4].
@@
Nonetheless, estimating resources often demands that end-users manually furnish extensive details about specific hardware characteristics, such as gate times and fidelities. Unlike previous research, our work does not require users to provide excessive hardware characteristic data or to have expert-level knowledge of quantum computing. Once a user writes a quantum program and selects a backend (either a simulator or a quantum computer) for execution, our methodology can automatically and accurately estimate the execution time of the quantum program on the chosen backend, making it highly accessible, efficient, and accurate. Another major difference is that our approach targets current backends (i.e., the simulators provided by Qiskit and quantum computers on the IBM Quantum Platform). This focus makes our model more practical and relevant for real-world applications.
2.2 Quantum program fidelity estimation
@@
Fidelity is a crucial measure that describes the closeness of a quantum state or operation to a reference state or operation, quantifying the degree of similarity between quantum states or the precision in quantum information preservation [5]. This measure is vital in practical quantum computing, where operations often face errors due to factors such as decoherence and imperfect gate implementations. High fidelity indicates minimal errors in quantum computations, crucial for reliable quantum computing. Recent strategies employ machine learning and deep learning to predict fidelity based on quantum circuit characteristics such as depth, and the type and number of quantum gates [29]. Additionally, machine learning classification methods are used to select the most suitable backend for compiling a specific quantum circuit to maximize fidelity [17, 20].
@@
2.3 Classical software execution time estimation
Current research has created methodologies that link the execution times of computer programs to their code features, enabling accurate predictions of execution times for conventional programs. Huang et al. [7] presented the SPORE (Sparse Polynomial Regression) methodology for predicting program execution time using a sparse, non-linear model based on features extracted from program execution data. Kwon et al. [10] developed Mantis, a framework designed to predict the execution time of Android applications. By leveraging program analysis and machine learning, Mantis identifies key program features strongly correlated with computational resource consumption (CRC) and automatically generates executable code snippets to evaluate them. Meng et al. [13] delved into an innovative nonlinear approach to estimate the worst-case execution time (WCET) during the programming phase. This approach uses historical WCET data from programs with similar object codes, including variations of those programs, to make predictions. However, practical challenges such as restricted access to source code due to licensing restrictions often emerge. Bielecki et al. [2] developed models that use program metadata and runtime environment parameters as explanatory variables, yet treat the program code itself as a black box. These models primarily aim to estimate the average-case execution time (ACET).

### v2

# Understanding and Estimating the Execution Time of Quantum Circuits

@@ Abstract.
Due to the scarcity of quantum computing resources, researchers and developers have very limited access to real quantum computers. Therefore, judicious planning and utilization of quantum computer runtime are essential to ensure smooth execution and completion of projects. Accurate estimation of a quantum circuit’s execution time is thus necessary to prevent unexpectedly exceeding the anticipated runtime or the maximum capacity of the quantum computers; it also allows quantum computing platforms to make precisely informed provisioning and prioritization of quantum computing jobs. In this paper, we first study the characteristics of quantum circuits’ runtime on simulators and real quantum computers. Then, we introduce an innovative method that employs a graph transformer-based model, utilizing the graph information and global information of quantum circuits to estimate their execution time. We selected a benchmark dataset comprising over 1,510 quantum circuits, initially predicting their execution times on simulators, which yielded promising results with an R-squared value greater than 95%. Subsequently, we applied active learning to select 340 circuit samples with a confidence level of 95% to build and evaluate our approach for the estimation of circuit execution times on quantum computers, achieving an average R-squared value exceeding 90%. Our approach can be integrated into quantum computing platforms to provide an accurate estimation of quantum execution time and be used as a reference for prioritizing quantum execution jobs. In addition, our findings provide insights for quantum program developers to optimize their circuits for reduced execution time.
@@
Quantum Computing, Quantum Circuits, Quantum Computers, Quantum Execution Time, Graph Transformer
journal: TOSEMccs: Software and its engineering Software performanceccs: Hardware Quantum computation
1. Introduction
@@
Over the last decade, there has been a significant transformation in quantum computing, evolving from experimental physics to cloud-accessible hardware (e.g., the IBM Quantum or Amazon Braket platforms), indicating a shift towards early-stage commercialization. This evolution has engaged a diverse audience, leading to research into practical quantum algorithms and gaining attention from academia, governments, and businesses. Notably, recent demonstrations of quantum utility (lundberg2017unified) suggest that quantum computers may eventually achieve comparable or better performance than classical counterparts for certain real-world tasks (zhong2020quantum).
@@
In quantum computing, the process of writing quantum programs111”Quantum programs” refers exclusively to circuit-level quantum programs in this paper. and making a quantum circuit executable on a specific backend is akin to classical computing’s transformation of high-level code into low-level, machine-specific code. Designed initially without considering hardware constraints such as gate limitations or connectivity issues, a quantum circuit undergoes a series of compilation passes (Quetschlich2023mqtbench). These passes progressively adapt the circuit to the target backend’s restrictions. Given the complexity of these compilation problems, there is a variety of available approaches in various quantum SDKs and tools. These SDKs differ in the programming languages, gate sets, and quantum computer vendors they support (salm2021automating). For example, vendor-agnostic SDKs such as pytket (sivarajah2020t) accommodate various languages and gate sets, whereas vendor-specific SDKs like IBM’s Qiskit (javadi2024quantum) are restricted to their proprietary systems.
@@
Executing a quantum circuit on a quantum computer involves exclusively using the quantum computer for a specific period, similar to the batch processing mode used in old mainframe systems. Users submit their designed and compiled quantum circuit to a pre-selected quantum computer, wait in a queue for processing, and finally the quantum computer executes the quantum circuit. The time taken to execute the quantum circuit is referred to as the ”execution time”. Given the scarcity and the substantial cost associated with quantum computing resources, the window available for using these state-of-the-art machines is exceptionally limited. Using IBM’s Quantum Platform [^2] https://quantum.ibm.com/ as an example, free users are limited to accessing two to three quantum computers per month, and usage is restricted to no more than 10 minutes. Any additional usage time is billed by the second and can be quite costly (\$1.6 USD per second at the time of writing). Exceeding time constraints can lead to incomplete executions or cause significant costs, potentially delaying critical projects or incurring other severe consequences. Prudent planning and strategic utilization of quantum computer time are imperative to ensure the smooth execution and completion of projects. This underscores the necessity for an accurate estimation of quantum circuits’ execution time. In addition, an accurate estimation of the execution time can help quantum computing platform providers (e.g., IBM Quantum Platform) with their provisioning and prioritization of quantum computing jobs.
@@
Despite the crucial need for reliable execution time predictions for quantum circuits, current research in this area remains sparse. Based on our experience of using the IBM Quantum Platform, its provided execution time estimate is often inaccurate (more details are discussed in Section 6), failing to meet the demands of practitioners and researchers in the quantum computing field adequately. There are existing studies on estimating the execution time of classical programs (huang2010predicting; kwon2014mantis; meng2016nonlinear; bielecki2023estimation). For instance, huang2010predicting collected a set of program execution features of a classical program, such as loop counts (the number of times a particular loop executed) and branch counts (the number of times a particular conditional branch executed), and build polynomial regression models to predict its execution time. This method achieved an accuracy of over 93% in predicting the execution time for the programs they tested. However, these approaches for classical programs do not apply to quantum programs, for three reasons: 1) quantum programs and classical programs are different in nature: classical programs describe the execution flows of code statements, whereas quantum programs describe the structure of a quantum circuit; 2) most approaches (huang2010predicting; kwon2014mantis; meng2016nonlinear) require tracing runtime information from the subject programs to collect features for model training or inference, whereas tracing quantum programs is not possible, at least for now, due to the famous no-cloning theorem (wootters1982single); 3) these approaches build a model for each program (huang2010predicting; kwon2014mantis; bielecki2023estimation) or similar programs (meng2016nonlinear) as different programs have very different features (e.g., different loops or branches), but it is too expensive to collect sufficient training data for each quantum program or similar ones, as quantum computing resources are rather limited.
@@
To address this gap, this paper introduces an innovative methodology that harnesses graph transformer-based models to predict the execution times of quantum circuits, leveraging the topological graph data intrinsic to these circuits and their global characteristics. We have analyzed over 1,510 quantum circuits, ranging in size from 2 to 127 qubits, and executed them using different backends: two simulators and two real quantum computers. We organize our study by answering three research questions.
@@
    Understanding the execution times of quantum circuits is crucial for efficiently managing quantum computing resources. Intuitively, different quantum circuits with varying structures might be expected to have different execution times. However, execution times are influenced by multiple complex factors, including the number of qubits, gate counts, types of quantum gates, and the specific characteristics of quantum backends, etc. Due to these interacting factors, it might be possible that quantum circuits with distinct structural features could exhibit similar execution times. If the variations in execution times across different quantum circuits on different quantum backends are minimal, detailed execution-time estimation might lose its practical importance. Therefore, this research question aims to verify whether the differences in execution times among various quantum circuits are sufficiently substantial. We observe that the execution times of different quantum circuits are noticeably different, whether on simulators or real quantum computers. In addition to these distributional differences, simple target-independent features (depth, num_qubits, two-qubit gate count, and num_gates) show nontrivial yet limited associations with execution time, and the associations differ across backends, indicating that single features are insufficient and motivating learned models that use backend-dependent features. Furthermore, the IBM Quantum Platform’s estimation of execution times for quantum circuits on quantum computers is not sufficient, which leads us to the second and third research questions.
@@
    Quantum program developers typically use simulators to verify their programs before executing them on real quantum computers. Time-consuming simulation can cost significant computing resources (e.g., IBM Quantum’s cloud-based simulators) and impact the productivity of quantum program development. Thus, understanding the simulation time of quantum circuits can help quantum developers better plan their simulation jobs (e.g., determining running a circuit for more or fewer shots) and help simulation service providers optimize the management of the simulation jobs. In this RQ, we utilize various features, including features that capture the overall characteristics of a circuit (i.e., global features and features that capture the graph structure of a circuit (i.e., graph features), and construct machine learning models to estimate the execution time of quantum circuits on simulators. Our results demonstrate the precision of our automated estimation and highlight the influential features.
@@
    Quantum computing resources are very limited currently and in the foreseeable future. Understanding the resource requirement (in terms of execution time) of quantum circuits before they execute is important for the optimal planning and allocation of quantum computing resources. Thus, this research question investigates the feasibility of creating a model that can accurately predict the execution times of quantum circuits on real quantum computers. Facing the limited quantum resources, we use active learning to collect the most representative training data. Our results show that we can achieve accurate predictions with a relatively small amount of data, substantially outperforming the estimation provided by the IBM Quantum Platform, and highlight the difference between the important predictors for the simulation time prediction and the prediction for real computer execution time.
@@
  • We provided an in-depth understanding of the execution time of quantum circuits on both simulators and real quantum computers.
@@
  • We identified the influential features that explain the quantum circuit execution time, providing insights for quantum program developers to consider when developing efficient quantum circuits.
@@
Paper organization. Section 2 discussed prior work related to our study. Section 3 introduces the background knowledge of quantum computing, including quantum circuits, compilation, and execution. Section 4 describes the data collection and preparation process. Section 5 presents our model construction and the evaluation methods. Section 6 presents and discusses our results for answering the research questions. Section 7 gives a further discussion of our experiment. Section 8 discusses the limitations of our study. Finally, Section 9 concludes our study and suggests directions for future research.
2. Related Works
@@
As far as we know, there exist no prior work on understanding or estimating the execution time of quantum circuits. In this section, we discuss three categories of prior works that are closest to ours: quantum resource estimation, quantum circuit fidelity estimation, and classical software execution time estimation.
2.1. Quantum resource estimation
@@
Quantum resource estimation offers a forward-looking solution for implementing key quantum algorithms by predicting necessary resources—such as qubits and runtime—based on future hardware projections, without running the actual programs. This approach allows consideration of quantum programs beyond current simulator and computer capabilities (quetschlich2024utilizingresourceestimationdevelopment). QuRE (6657074) was among the pioneering efforts on quantum resource estimation. The authors developed a resource estimation methodology for diverse quantum technologies and algorithms, capable of automatically estimating metrics like qubit count, completion time, success probability, and gate counts for different physical gate types. Later on, another work has been undertaken to develop a framework for quantum resource estimation and evaluate the resource requirements for large-scale quantum applications, with a particular emphasis on the critical role of qubit parameters, aiming to expedite progress towards achieving practical quantum advantage (beverland2022assessing). Another recent research provided resource estimates for quantum derivative pricing; it introduced a resource-efficient re-parameterization method, offered insights into resource demands for benchmark use cases, and outlined a roadmap for future quantum algorithm and hardware enhancements (Chakrabarti2021thresholdquantum).
Although tools such as QuRE have been developed to estimate quantum execution time, their applicable scenario is fundamentally different from our work. QuRE is designed as a theoretical resource estimation framework for fault-tolerant quantum computing. It estimates the total time to completion by modeling full-stack implementations of quantum algorithms, incorporating fault-tolerant error correction (e.g., Steane or surface codes), magic state distillation, tile-based layouts, and low-level physical gate scheduling. These estimates are based on hypothetical hardware models and assume the use of logical qubits constructed from many physical qubits, often requiring millions to billions of gates. In contrast, our work focuses on real-world near-term quantum computing, targeting NISQ devices. The execution time we estimate corresponds to the actual latency experienced by quantum circuits on realistic quantum backends, after compilation and transpilation. Our metric captures compilation-specific characteristics such as qubit mapping, SWAP insertion, and hardware-level gate duration. Thus, while both works use the term ”execution time”, the underlying definitions and contexts are not directly comparable.
@@
Estimating resources often demands that end-users manually furnish extensive details about specific hardware characteristics, such as gate times and fidelities. Unlike previous research, our work does not require users to provide excessive hardware characteristic data or to have expert-level knowledge of quantum computing. Once a user writes a quantum program and selects a backend (either a simulator or a quantum computer) for execution, our methodology can automatically and accurately estimate the execution time of the quantum circuit on the chosen backend, making it highly accessible, efficient, and accurate. Another major difference is that our approach targets current backends (i.e., the simulators provided by Qiskit and quantum computers on the IBM Quantum Platform). This focus makes our model more practical and relevant for real-world applications.
2.2. Quantum circuit fidelity estimation
@@
Fidelity is a crucial measure that describes the closeness of a quantum state or operation to a reference state or operation, quantifying the degree of similarity between quantum states or the precision in quantum information preservation (gilyen2022improved). This measure is vital in practical quantum computing, where operations often face errors due to factors such as decoherence and imperfect gate implementations. High fidelity indicates minimal errors in quantum computations, crucial for reliable quantum computing. Recent strategies employ machine learning and deep learning to predict fidelity based on quantum circuit characteristics such as depth, and the type and number of quantum gates (vadali2024quantum). Additionally, machine learning classification methods are used to select the most suitable backend for compiling a specific quantum circuit to maximize fidelity (quetschlich2023compiler; quetschlich2023predicting).
@@
2.3. Classical software execution time estimation
Current research has created methodologies that link the execution times of computer programs to their code features, enabling accurate predictions of execution times for conventional programs.  huang2010predicting presents the SPORE (Sparse Polynomial Regression) methodology for predicting program execution time using a sparse, non-linear model based on features extracted from program execution data. kwon2014mantis developed Mantis, a framework designed to predict the execution time of Android applications. By leveraging program analysis and machine learning, Mantis identifies key program features strongly correlated with computational resource consumption (CRC) and automatically generates executable code snippets to evaluate them. meng2016nonlinear delves into an innovative nonlinear approach to estimate the worst-case execution time (WCET) during the programming phase. This approach uses historical WCET data from programs with similar object codes, including variations of those programs, to make predictions. However, practical challenges such as restricted access to source code due to licensing restrictions often emerge. bielecki2023estimation develop models that use program metadata and runtime environment parameters as explanatory variables, yet treat the program code itself as a black box. These models primarily aim to estimate the average-case execution time (ACET).